# New Applications of Ergodic Theory to Sets of Differences

### v1

@@ 5 Examples of tempered Følner sequences
In this section we collect some examples of tempered Følner sequences. In Subsection 5.1 we present two types of tempered sequences in $(\mathbb{N},\cdot)$. In Subsection 5.2 we provide an example of a tempered Følner sequence in the Heisenberg group. Finally, in Subsection 5.3 we observe that the natural Følner sequences in locally finite groups are actually tempered.
@@ 5.1 Tempered Sequences in $(\mathbb{N},\cdot)$
Proof: Let $g\in(\mathbb{N},\cdot)$ with prime factorization $g=p_{1}^{r_{1}}\dots p_{k}^{r_{k}}$ where $r_{j}\geq 0$ for $1\leq j\leq k$. Then for all $n>k$,
$F_{n}\cap gF_{n}=\left\{p_{1}^{\ell_{1}}\dots p_{n}^{\ell_{n}}:r_{i}\leq\ell_{i}\leq(i+1)^{2n}\text{ for }1\leq i\leq k,\text{ }0\leq\ell_{i}\leq(i+1)^{2n}\text{ for }i>k\right\}.$
This means that
@@
Also note that
$|F_{n}|=\prod_{i=1}^{n}((i+1)^{2n}+1).$
We have
$\lim_{n\to\infty}\frac{|F_{n}\cap gF_{n}|}{|F_{n}|}=\lim_{n\to\infty}\prod_{i=1}^{k}\frac{(i+1)^{2n}+1-r_{i}}{(i+1)^{2n}+1}=1.$
Thus $(F_{n})_{n=1}^{\infty}$ is a Følner sequence. If we show $(F_{n})_{n=1}^{\infty}$ is tempered in $\mathbb{Q}^{\times}_{>0}$ then Theorem 4.3 will imply $(F_{n})_{n=1}^{\infty}$ is tempered in $(\mathbb{N},\cdot)$. For this reason we will regard $(F_{n})_{n=1}^{\infty}$ as a sequence of sets in $\mathbb{Q}_{>0}^{\times}$ for the remainder of the proof. Let $n>1$ and notice that $F_{k}\subseteq F_{n-1}$ for $11<k\leq n-$ 1, so $\bigcup_{k=1}^{n-1}F_{k}^{-1}F_{n}=F_{n-1}^{-1}F_{n}$. Also note that
$F_{n-1}^{-1}F_{n}=\left\{p_{1}^{\ell_{1}}\dots p_{n}^{\ell_{n}}:-(i+1)^{2(n-1)}\leq\ell_{i}\leq(i+1)^{2n}\text{ for }1\leq i\leq n-1,\hskip 5.69054pt0\leq\ell_{n}\leq(n+1)^{2n}\right\},$
Thus
$|F_{n-1}^{-1}F_{n}|=((n+1)^{2n}+1)\prod_{i=1}^{n-1}((i+1)^{2(n-1)}+(i+1)^{2n}+1).$
So
@@
Since $\prod_{i=1}^{\infty}\left(1+\frac{1}{(i+1)^{2}}\right)$ converges to some value $C>0$, we have that for all $n>1$,
@@
and hence $(F_{n})_{n=1}^{\infty}$ is a tempered Følner sequence in $(\mathbb{N},\cdot)$. ∎
Remark 5.2. Note that for any $\epsilon>0$ if we let, for $n\in\mathbb{N}$, $F_{n}=\{p_{1}^{c_{1}}\dots p_{n}^{c_{n}}:0\leq c_{i}\leq(i+1)^{(1+\epsilon)n}\}$, then it follows from the proof of Theorem 5.1 that $(F_{n})_{n=1}^{\infty}$ is tempered. Now let, for $n\in\mathbb{N}$, $G_{n}=\{p_{1}^{c_{1}}\dots p_{n}^{c_{n}}:0\leq c_{i}\leq(i+1)^{n}\}$. One can show that $(G_{n})_{n=1}^{\infty}$ is not tempered.
Theorem 5.3. Let $f:\mathbb{N}\rightarrow\mathbb{N}$ be a nondecreasing function such that $f(n)\rightarrow\infty$. For $n\in\mathbb{N}$ let
$F_{n}=\{p_{1}^{c_{1}}\cdots p_{N}^{c_{N}}:0\leq c_{i}\leq f(N)\text{ for }1\leq i\leq N\}.$
Then $(F_{n})_{n=1}^{\infty}$ is a tempered Følner sequence in $(\mathbb{N},\cdot)$ iff the sequence
$\frac{nf(n)}{f(n+1)}$ (5.1)
@@
Proof. By Theorem 4.3, it suffices to show $(F_{n})$ is a tempered sequence in $\mathbb{Q}_{>0}^{\times}$, so we will regard $(F_{n})$ as a sequence in $\mathbb{Q}^{\times}_{>0}$ for the remainder of the proof. Since $f$ is non-decreasing, $\left|\bigcup_{k=1}^{n-1}F_{k}^{-1}F_{n}\right|=|F_{n-1}^{-1}F_{n}|=(f(n)+1)(f(n)+f(n-1)+1)^{n-1}$ for $n\in\mathbb{N}$ with $n>1$, hence
$\frac{|\bigcup_{k=1}^{n-1}F_{k}^{-1}F_{n}|}{|F_{n}|}=\left(1+\frac{f(n-1)}{f(n)+1}\right)^{n-1}.$ (5.2)
It follows that $(F_{n})_{n=1}^{\infty}$ is tempered iff the sequence
$\left(1+\frac{f(n-1)}{f(n)+1}\right)^{n-1}$ (5.3)
is bounded. Since $f(n)\to\infty$, one can verify that $\frac{(n-1)f(n-1)}{f(n)+1}$ is bounded iff $\frac{nf(n)}{f(n+1)}$ is bounded, so it suffices to show that the term in (5.3) is bounded iff $\frac{(n-1)f(n-1)}{f(n)+1}$ is bounded.
()($\Rightarrow$)( ⇒ ) First suppose that $\frac{(n-1)f(n-1)}{f(n)+1}$ is not bounded. Let $C>0$. Choose $M\in\mathbb{N}$ such that $\frac{(M-1)f(M-1)}{f(M)+1}>C$. Then we have
$\left(1+\frac{f(M-1)}{f(M)+1}\right)^{M-1}$ $=\sum_{k=0}^{M-1}\binom{M-1}{k}\left(\frac{f(M-1)}{f(M)+1}\right)^{k}>\frac{(M-1)f(M-1)}{f(M)+1}>C.$ (5.4)
Thus for all $C>0$ there is some $M\in\mathbb{N}$ such that
$\left(1+\frac{f(M-1)}{f(M)+1}\right)^{M-1}>C,$ (5.5)
so the term in (5.3) is unbounded.
($\Leftarrow$) Now suppose that $\frac{(n-1)f(n-1)}{(f(n)+1)}$ is bounded above by $C>0$. Then since $f$ is non-decreasing for all $n\in\mathbb{N}$ we have
$\frac{nf(n-1)}{f(n)+1}<C+1.$
Note that for $n\in\mathbb{N}$ with $n>1$ and for all $1<k\leq n$ we have $n!/(n-k)!<n^{k}$, and so
$\left(1+\frac{f(n-1)}{f(n)+1}\right)^{n-1}<\left(1+\frac{f(n-1)}{f(n)+1}\right)^{n}$ $=\sum_{k=0}^{n}\binom{n}{k}\left(\frac{f(n-1)}{f(n)+1}\right)^{k}$ (5.6)
$=\sum_{k=0}^{n}\frac{n!}{k!(n-k)!}\left(\frac{f(n-1)}{f(n)+1}\right)^{k}$ (5.7)
$<\sum_{k=0}^{n}\frac{1}{k!}\left(\frac{nf(n-1)}{f(n)+1}\right)^{k}<\sum_{k=0}^{\infty}\frac{(C+1)^{k}}{k!}=e^{C+1}.$ (5.8)
Thus the term in (5.3) is bounded above by $e^{C+1}$. ∎
Remark 5.4. Let $k\in\mathbb{N}$, for $n\in\mathbb{N}$ define
$F_{n}=\{p_{1}^{c_{1}}\cdots p_{n}^{c_{n}}:0\leq c_{1},\ldots,c_{n}\leq n^{k}\}.$
@@
is a tempered Følner sequence in $(\mathbb{N},\cdot)$
@@ 5.2 Tempered Sequences in the Heisenberg group
Proof.
In what follows we will write $(a,b,c)$ to denote the matrix
@@
Then for $n\in\mathbb{N}$ we have
$F_{n}=\{(a,b,c):a,b\in\{-n,\ldots,n\}\text{ and }c\in\{-n^{2},\ldots,n^{2}\}\}.$
It is straightforward, albeit tedious, to verify that $(F_{n})_{n=1}^{\infty}$ is a Følner sequence in $H$.
For $(x,y,z)\in H$ we have $(x,y,z)^{-1}=(-x,-y,xy-z)$. Then for $n\in\mathbb{N}$ with $n>1$ we see that
$F_{n-1}^{-1}=\{(a,b,c):-(n-1)\leq a,b\leq n-1\text{ and }-(n-1)^{2}+ab\leq c\leq(n-1)^{2}+ab\},$
so
@@
$\leq(2(2n-1)+1)^{2}(2(n-1)^{2}+2n^{2}+2n(n-1)+2n^{2}+1).$
It follows that
$\frac{|\bigcup_{k=1}^{n-1}F_{k}^{-1}F_{n}|}{|F_{n}|}\leq\frac{(2(2n-1)+1)^{2}(2(n-1)^{2}+2n^{2}+2n(n-1)+2n^{2}+1)}{(2n+1)^{2}(2n^{2}+1)}\rightarrow\text{constant},$
so $(F_{n})_{n=1}^{\infty}$ is left tempered. ∎
Remark 5.6. A similar argument shows that sets $(F_{n})_{n=1}^{\infty}$ defined by $F_{n}=\{(a,b,c):a,b\in\{-n,\ldots,n\}\text{ and }c\in\{-p(n),\ldots,p(n)\}\}$ form a left tempered Følner sequence in the Heisenberg group $H$ whenever $p(x)\in\mathbb{Z}[x]$ is a positive polynomial of degree at least 2.

### v2

@@ 5 Examples of tempered Følner sequences
In this section we collect some examples of tempered Følner sequences, which, as we saw in subsection 4.2, play an important role in extending Theorem 1 to the framework of amenable groups. In subsection 5.1 we present some useful facts and examples relating to tempered and non-tempered Følner sequences in $(\mathbb{N},+)$. In subsection 5.2 we provide examples of tempered Følner sequences in $(\mathbb{N},\cdot)$. In subsection 5.3 we provide an example of a tempered Følner sequence in the Heisenberg group. Finally, in subsection 5.4 we observe that the natural Følner sequences in locally finite groups are actually tempered.
5.1 Tempered Sequences in $(\mathbb{N},+)$
It is well-known that the sequence $(F_{N})_{N=1}^{\infty}$, where $F_{N}=\{1,\ldots,N\}$, is a tempered Følner sequence in $(\mathbb{N},+)$. The following simple observation is a generalization of this fact:
Theorem 5.1. For $a,b\in\mathbb{N}$, $a\leq b$, let $[a,b]$ denote the set $\{a,a+1,\dots,b-1,b\}$. If $(a_{n})_{n=1}^{\infty}$ and $(b_{n})_{n=1}^{\infty}$ are sequences in $\mathbb{N}$ such that $b_{n}-a_{n}\to\infty$ and $[a_{j},b_{j}]\subseteq[a_{j+1},b_{j+1}]$ for all $j\in\mathbb{N}$, then $F_{n}:=[a_{n},b_{n}]$, $n=1,2,3,...$ is a tempered Følner sequence in $\mathbb{N}$.
Proof. In view of Theorem 4.3, it suffices to show that $(F_{n})$ is a tempered Følner sequence in $(\mathbb{Z},+)$. For this reason, any differences $F_{i}-F_{j}$ will be computed in $(\mathbb{Z},+)$. [^11] Given $A,B\subseteq\mathbb{Z}$, we define $A-B=\left\{a-b:a\in A,b\in B\right\}.$ Note that this is the additive analog of formula (4.15). It is clear that $(F_{n})$ is a Følner sequence in $(\mathbb{Z},+)$. Since $F_{1}\subseteq F_{2}\subseteq F_{3}\subseteq\cdots$, it follows that for any $n\in\mathbb{N}$ with $n>1$, we have
$\left|\bigcup_{1\leq k<n}(F_{n}-F_{k})\right|$ $=|F_{n}-F_{n-1}|=\Big{|}[a_{n}-b_{n-1},b_{n}-a_{n-1}]\Big{|}$
$=b_{n}-a_{n-1}+b_{n-1}-a_{n}+1\leq 2b_{n}-2a_{n}+1<2(b_{n}-a_{n}+1)=2|F_{n}|.$
This shows that $(F_{n})$ is tempered. $\square$
In [1] it was shown that the pointwise ergodic theorem fails along the sequence $([n,n+[\sqrt{n}]])_{n=1}^{\infty}$, so it follows from Theorem 4.2 that the Følner sequence $([n,n+[\sqrt{n}]])_{n=1}^{\infty}$ is not tempered. This result was later strengthened in [17, Theorem 1] and [16, Corollary 3.9], where it was shown that for any nondecreasing sequence $(b_{n})_{n=1}^{\infty}$ in $\mathbb{N}$ such that $\lim_{n\to\infty}n/b_{n}=0$, the pointwise ergodic theorem fails along the Følner sequence $([n,n+b_{n}])_{n=1}^{\infty}$, and hence $([n,n+b_{n}])_{n=1}^{\infty}$ is not tempered. Another example of a Følner sequence that is not tempered is the sequence $([n^{2},n^{2}+n])_{n=1}^{\infty}$ (see [11, Corollary 6] and also [25, Example 2.7]). Since every Følner sequence has a tempered subsequence, it is natural to ask for an example of a tempered subsequence of $([n^{2},n^{2}+n])_{n=1}^{\infty}$ that is tempered. One such example was provided in [11, page 53], where it was shown that $([2^{2^{n}},2^{2^{n}}+2^{2^{n-1}}])_{n=1}^{\infty}$ is tempered subsequence of $([n^{2},n^{2}+n])_{n=1}^{\infty}$.
@@ 5.2 Tempered Sequences in $(\mathbb{N},\cdot)$
Proof. We first show that $(F_{n})$ is a Følner sequence. To this end, let $g\in\mathbb{N}$ have prime factorization $g=p_{1}^{r_{1}}\dots p_{k}^{r_{k}}$, where $r_{j}\geq 0$ for $1\leq j\leq k$, and observe that for all $n>k$,
$F_{n}\cap gF_{n}=\left\{p_{1}^{\ell_{1}}\dots p_{n}^{\ell_{n}}:r_{i}\leq\ell_{i}\leq(i+1)^{2n}\text{ for }1\leq i\leq k,\text{ and }0\leq\ell_{i}\leq(i+1)^{2n}\text{ for }i>k\right\}.$
This means that, for sufficiently large $n$,
@@
Since $|F_{n}|=\prod_{i=1}^{n}((i+1)^{2n}+1)$ for all $n\in\mathbb{N}$, we have
$\lim_{n\to\infty}\frac{|F_{n}\cap gF_{n}|}{|F_{n}|}=\lim_{n\to\infty}\prod_{i=1}^{k}\frac{(i+1)^{2n}+1-r_{i}}{(i+1)^{2n}+1}=1,$
and so $(F_{n})_{n=1}^{\infty}$ is a Følner sequence in $(\mathbb{N},\cdot)$. It follows from Theorem 4.3 that $(F_{n})_{n=1}^{\infty}$ is also a Følner sequence in $\mathbb{Q}_{>0}^{\times}$. Thus, by Theorem 4.3, it suffices to show that $(F_{n})_{n=1}^{\infty}$ is tempered in $\mathbb{Q}_{>0}^{\times}$. For this reason, the remainder of our calculations will be carried out in the group $\mathbb{Q}_{>0}^{\times}$. Let $n\in\mathbb{N}$ with $n>1$ and observe that since $F_{1}\subseteq F_{2}\subseteq F_{3}\subseteq\cdots$, we have
$\bigcup_{k=1}^{n-1}F_{k}^{-1}$ $F_{n}=F_{n-1}^{-1}F_{n}$
[…]
Therefore,
@@
Let $C=\prod_{i=1}^{\infty}\left(1+\frac{1}{(i+1)^{2}}\right)$. Notice $C<\infty$. We have that for all $n>1$,
@@
and hence $(F_{n})_{n=1}^{\infty}$ is a tempered Følner sequence in $(\mathbb{N},\cdot)$. $\square$
Remark 5.3. Note that for any $\epsilon>0$ if we let $F_{n}=\{p_{1}^{c_{1}}\dots p_{n}^{c_{n}}:0\leq c_{i}\leq(i+1)^{(1+\epsilon)n}\}$ for all $n\in\mathbb{N}$, then an argument similar to the proof of Theorem 5.2 shows that $(F_{n})_{n=1}^{\infty}$ is tempered. On the other hand, if $G_{n}=\{p_{1}^{c_{1}}\dots p_{n}^{c_{n}}:0\leq c_{i}\leq(i+1)^{n}\}$ for all $n\in\mathbb{N}$, one can show that $(G_{n})_{n=1}^{\infty}$ is not tempered.
Theorem 5.4. Let $f:\mathbb{N}\rightarrow\mathbb{N}$ be a nondecreasing function such that $f(n)\rightarrow\infty$. For $n\in\mathbb{N}$ let $F_{n}=\{p_{1}^{c_{1}}\cdots p_{n}^{c_{n}}:0\leq c_{i}\leq f(n)\text{ for }1\leq i\leq n\}$. Then $(F_{n})_{n=1}^{\infty}$ is a Følner sequence in $(\mathbb{N},\cdot)$ and it is tempered iff the sequence
$\frac{nf(n)}{f(n+1)},\quad n\in\mathbb{N}$ (5.1)
@@
Proof. It is straightforward to verify that $(F_{n})$ is a Følner sequence in $(\mathbb{N},\cdot)$, so we only show that $(F_{n})$ is tempered iff the sequence in formula (5.1) is bounded. By Theorem 4.3, it suffices to show that $(F_{n})$ is a tempered Følner sequence in $\mathbb{Q}_{>0}^{\times}$ iff the sequence in formula (5.1) is bounded, so we will carry out all of our calculations in $\mathbb{Q}^{\times}_{>0}$ for the remainder of the proof. Since $f$ is non-decreasing, for all $n\in\mathbb{N}$ with $n>1$, we have
$\frac{\left|\bigcup_{k=1}^{n-1}F_{k}^{-1}F_{n}\right|}{\left|F_{n}\right|}$ $=\frac{\left|F_{n-1}^{-1}F_{n}\right|}{\left|F_{n}\right|}=\frac{\left|\left\{p_{1}^{\ell_{1}}\cdots p_{n}^{\ell_{n}}:-f(n-1)\leq\ell_{1},\ell_{2},...,\ell_{n-1}\leq f(n),0\leq\ell_{n}\leq f(n)\right\}\right|}{\left|F_{n}\right|}$
$=\frac{(f(n)+1)(f(n)+f(n-1)+1)^{n-1}}{(f(n)+1)^{n}}=\left(1+\frac{f(n-1)}{f(n)+1}\right)^{n-1}.$
Hence, by the definition of temperedness, $(F_{n})_{n=1}^{\infty}$ is tempered iff the sequence
$a_{n}:=\left(1+\frac{f(n-1)}{f(n)+1}\right)^{n-1},\quad n\in\mathbb{N}$ (5.2)
is bounded.
Now note that the sequence in formula (5.1) is bounded iff the sequence $b_{n}:=\frac{(n-1)f(n-1)}{f(n)+1}$, $n\in\mathbb{N}$ is bounded. Thus, it suffices to show that $(a_{n})_{n=1}^{\infty}$ is bounded iff the $(b_{n})_{n=1}^{\infty}$ is bounded.
()($\Rightarrow$)( ⇒ ) Suppose that $(a_{n})_{n=1}^{\infty}$ is bounded. Since for every $n>1$, we have
$b_{n}=\frac{(n-1)f(n-1)}{f(n)+1}\leq\sum_{k=0}^{n-1}\binom{n-1}{k}\left(\frac{f(n-1)}{f(n)+1}\right)^{k}<\left(1+\frac{f(n-1)}{f(n)+1}\right)^{n-1}=a_{n},$
it follows that $(b_{n})_{n=1}^{\infty}$ is bounded as well.
()($\Leftarrow$)( ⇐ ) Suppose that $(b_{n})_{n=1}^{\infty}$ is bounded above by some $C>0$. Since $f$ is non-decreasing, for all $n\in\mathbb{N}$, we have $\frac{nf(n-1)}{f(n)+1}\leq\frac{(n-1)f(n-1)}{f(n)+1}+1=b_{n}+1<C+1$. It follows that for any $n\in\mathbb{N}$ with $n>1$, we have
$a_{n}=\left(1+\frac{f(n-1)}{f(n)+1}\right)^{n-1}<\left(1+\frac{f(n-1)}{f(n)+1}\right)^{n}=\sum_{k=0}^{n}\frac{n(n-1)\cdots(n-k+1)}{k!}\left(\frac{f(n-1)}{f(n)+1}\right)^{k}$
$<\sum_{k=0}^{n}\frac{n^{k}}{k!}\left(\frac{f(n-1)}{f(n)+1}\right)^{k}=\sum_{k=0}^{n}\frac{1}{k!}\left(\frac{nf(n-1)}{f(n)+1}\right)^{k}<\sum_{k=0}^{n}\frac{(C+1)^{k}}{k!}<\sum_{k=0}^{\infty}\frac{(C+1)^{k}}{k!}=e^{C+1}.$
Thus, $(a_{n})_{n=1}^{\infty}$ is bounded above by $e^{C+1}$. $\square$
Remark 5.5. Let $k\in\mathbb{N}$, and define
$F_{n}=\{p_{1}^{c_{1}}\cdots p_{n}^{c_{n}}:0\leq c_{1},\ldots,c_{n}\leq n^{k}\},\quad n=1,2,3,...$
@@
is a tempered Følner sequence in $(\mathbb{N},\cdot)$.
@@ 5.3 Tempered Sequences in the Heisenberg group
Proof. In what follows, we will write $(a,b,c)$ to denote the matrix
@@
Then for $n\in\mathbb{N}$, we have $F_{n}=\{(a,b,c):a,b\in\{-n,\ldots,n\}\text{ and }c\in\{-n^{2},\ldots,n^{2}\}\}$. It is a well-known fact that $(F_{n})_{n=1}^{\infty}$ is a Følner sequence in $H$, so we proceed to proving that $(F_{n})_{n=1}^{\infty}$ is tempered. For $(x,y,z)\in H$, we have $(x,y,z)^{-1}=(-x,-y,xy-z)$, so for $n\in\mathbb{N}$ with $n>1$, we see that
$F_{n-1}^{-1}$ =$=\{(-x,-y,xy-z):-(n-1)\leq x,y\leq n-1\text{ and }-(n-1)^{2}\leq z\leq(n-1)^{2}\}$
=$=\{(x,y,z):-(n-1)\leq x,y\leq n-1\text{ and }-(n-1)^{2}+xy\leq z\leq(n-1)^{2}+xy\}.$
For $(x,y,z),(a,b,c)\in H$, one has $(x,y,z)\cdot(a,b,c)=(x+a,y+b,z+xb+c)$. Therefore, for $n>1$,
@@
|$\leq\left|\left\{x+a:\begin{array}[]{l}-(n-1)\leq x\leq n-1\\ -n\leq a\leq n\\ \end{array}\right\}\right|\cdot\left|\left\{y+b:\begin{array}[]{l}-(n-1)\leq y\leq n-1\\ -n\leq b\leq n\\ \end{array}\right\}\right|$
|$\hphantom{hi}\quad\quad\cdot\left|\left\{z+xb+c:\begin{array}[]{l}-(n-1)\leq x,y\leq n-1\\ -(n-1)^{2}+xy\leq z\leq(n-1)^{2}+xy\\ -n\leq b\leq n\\ -n^{2}\leq c\leq n^{2}\end{array}\right\}\right|$
$\leq(2(2n-1)+1)^{2}(4(n-1)^{2}+2n(n-1)+2n^{2}+1).$
It follows that, for all $n>1$,
$\frac{|\bigcup_{k<n}F_{k}^{-1}F_{n}|}{|F_{n}|}\leq\frac{(2(2n-1)+1)^{2}(4(n-1)^{2}+2n(n-1)+2n^{2}+1)}{(2n+1)^{2}(2n^{2}+1)}.$ (5.3)
Let $f(n)$ denote the expression on the right-hand side of formula (5.3). Note that $\lim_{n\to\infty}f(n)$ exists and is finite. Let $C=\sup_{n\in\mathbb{N}}f(n)$. It follows from formula (5.3) that $\left|\bigcup_{k<n}F_{k}^{-1}F_{n}\right|\leq C\left|F_{n}\right|$ for all $n\in\mathbb{N}$. Thus, $(F_{n})_{n=1}^{\infty}$ is tempered. $\square$
Remark 5.7. A similar argument shows that the sequence
$F_{n}=\left\{(a,b,c):a,b\in\left\{-p(n),...,p(n)\right\},c\in\left\{-q(n),...,q(n)\right\}\right\},\quad n=1,2,3,...$
forms a left tempered Følner sequence in the Heisenberg group $H$ whenever $p(x),q(x)\in\mathbb{Z}[x]$ are positive polynomials with $\deg(q)\geq 2\deg(p)$.